17. Find the indicated probability. Round to the nearest thousandth. (I point) A study conducted at a certain college shows that 55% of the schoors graduates find a job in their chosen field within a year after graduation. Find the probability that among 7 randomly selected graduates, at least one finds a job in his or her chosen field within a year of O 0.550 O 0.996 0985 00.143
18. Find the indicated probability. Express your answer as a simplified fraction unless otherwise noted. (lpoimty The following table contains data from a study of two airlines which fly to Small Town, USA. Number of flightsNumber of flights which were on time which were late Podunk Airlines Upstate Airlines If one of the 87 flights is randomly selected, find the probability that the flight selected is arn 6 43 Upstate Airlines flight which was on time. 76 43 87 43 76 43 48
19. A company wanted to determine what percentage of its employees stayed for at least 1, 2, and (I point 3 years. The data they compiled is given below and covers 200 male employees and 300 female employees. 2 years20 Male Female 0.67 0.73 0.45 0.64 0.39 What is the probability that an employee stayed for at least one year? Round your answer to the nearest hundredth. o 0.70 o 0.71 d 0.06 140
20. Find the indicated probability. Express your answer as a simplified fraction unless (1 point) otherwise noted. The following table contains data from a study of two airlines which fly to Small Town, USA Number of flightsNumber of flights which were on time which were late Podunk Airlines Upstate Airlines 6 43 If one of the 87 flights is randomly selected, find the probability that the flight selected arrived on time given that it was an Upstate Airlines flight. 76 43 87 43 48 43 5
21. A pen is dropped on a hardwood floor and a 1 is recorded if the pen crosses a line and a 0ifit point) does not. Estimate the probability that the pen crosses a line in a single drop. 1110 10010011 1 8 13 13 13
22. A light can shine blue (B), red (R), or yellow (Y). Every second the color changes and is( point) recorded. Estimate the probability the next color will be red. BRYRRRRYRBRR 2 3 2
23. On his way to work each day, Steve encounters one traffic light. He records the color of the (lpoint) light each day by marking G for green and R for red. Estimate the probability that Steve encounters a green light on any given day GGGRRGRGGRRRRGGGGRRG 20 2 20 o 20 2 20
24. For every truck a police officer encounters she writes down T, for every car C, for every van V, (l point) and for every vehicle that is not a car, truck or van O (for other). Estimate the probability that the next vehicle the police officer encounters is not a car, truck, or van TTTC VOOVCTTCcCCCCVO 3 19 16 19 13 19 19
Answer #1

17)

P(at least one find am job)=1-P(none find a job)=1-(1-0.55)7 =0.996

18)

P(upstate airline and on time)=43/87

19)

P(at least 1 year)=P(male)*P(at least 1 year|male)+P(female)*P(at least 1 year|female)

=(2/5)*0.67+(3/5)*0.73=0.71

20)

P(on time |upstate)=43/48

21)

8/13

22)

8/12=2/3

23)

11/20

24)

3/19